OpenStudy (anonymous):

Calculate the sale price of this item before taxes. 30% off a bike for $397.95

OpenStudy (lilg132):

397.95 / 100 = 3.9795 (this is 1% of the price) 3.9795 * 30 = 119.385 = 119.39 (this is 30% of the price) so you do this 397.95 - 119.39 = $278.56 is the total price with 30% off dont forget to click good answer :)